steel that is 5% nickel and 95% iron is mixed with another batch that is 2% nickel and 98% iron to make 3 tons of steel that is 4% nickel and 96% iron. how many tons of each original batch was used?
I'm confused of which formula to use....

Answer by Boreal(15235)

x=5% batch in tons
3-x=2% batch
x(.05)+(3-x)(.02)=3(.04)
.05x + .06-.02x= 0.12
0.03x=0.06
x=2 tons of 5%
3-x=1 ton of 2%